# Specific heat and density anomaly in the Hubbard model

### v1

@@ III Results
For the Hubbard model, it has been established that this approximation to the complete Seebeck coefficient is one of the best available [52], even being in excellent quantitative agreement with experiments for some materials [54]. Because the denominator is always positive, a sign change of the Seebeck coefficient must be induced by a sign change of $\alpha_{\mu}$. Namely, the reported Seebeck anomaly [22] regarding its sign, and hence carrier charge, is controlled by the thermal expansion, with its magnitude modulated by the isothermal compressibility. This connection can be conceptually visualized as follows: a given temperature gradient $\delta T>0$ will induce a density change $\delta n>0$, as long as $\alpha_{\mu}$ is positive. This charge variation will induce an electric field in the same direction, and hence we get a positive Seebeck coefficient. On the other hand, an opposite response of density $\delta n<0$, due to a negative $\alpha_{\mu}$, will establish an electric field in the opposite direction, yielding a negative Seebeck response. Near half-filling and for strong interactions, $n^{2}\kappa_{T}$ goes to zero [22], greatly enhancing the Seebeck coefficient. At half-filling, the compressibility goes to zero with decreasing temperature, but the thermal expansion coefficient is also zero, leading to a zero Seebeck coefficient. It also goes to zero at the extreme values of $n=0$ and $n=2$, generating the Seebeck divergences for the empty and completely filled lattices.
Figure 7: Entropy $s$ as a function of density $n$ for a) $T=2$, b) $T=1$ and c) $T=0.5$ considering interactions ranging from the weak to the strong coupling limit. The $U=0$ line is hidden behind the $U=2$ case.
Studies of classical lattice models connected a change of sign in $\alpha_{\mu}$ to residual entropies [16], as well as the results for bosons [14, 10]. Although the ground state of the Fermi-Hubbard model has no residual entropy in the thermodynamic limit, at finite temperature, the entropy as a function of lattice filling $s(n)$ has maxima away from half-filling for strong correlations [22]. From Equation 6, we can see that $\alpha_{\mu}=0$ implies $(\partial s/\partial n)_{T}=0$. That is, the change of sign of $\alpha_{\mu}$ happens right at the entropy maxima. Conversely, the same is true for the Seebeck coefficient: its anomalous change of sign happens at the maxima in the entropy. Figure 7 shows that as the temperature is decreased, the entropy decreases for all interaction strengths and lattice fillings, but we detect the known maxima away from half-filling for all temperatures. For lower temperatures, the maxima develop at weaker interactions: while for $T=2$ only $U=8$ and $U=10$ show maxima away from half-filling, at $T=0.5$ this effect begins at $U=4$ already. We note that these maxima are a clear sign of the Mott insulator state: at $n=1$ and strong interaction, the entropy is at a local minimum. Any doping drives the system to a metallic state, leading to a larger entropy. As the doping increases further away from half-filling, carrier mobility is hindered, reducing the entropy. At the extreme values of $n=0$ and $n=2$ the system is a band insulator, leading to a very low entropy. In other words, the entropic competition between the metal and the insulators at each value of $n$ leads to the non-monotonic behavior of the entropy. At the $n$ where the entropy is maximum, the winner of this competition changes. To map the location of these effects as temperature changes, we mark on Figure 8 the points where $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=\partial s/\partial n=0$, for all the simulated temperatures and interactions. For the higher temperature, strong interactions ($U>8$) are necessary for the change of sign to happen. As the temperature is lowered, the effect starts to occur at a lower value of $U$ and a wider range of interactions is included.
Figure 8: Location of $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=(\partial s/\partial n)_{T}=0$ for all the simulated temperatures and interactions, estimated by the average between densities with opposite $\alpha_{\mu}$ signs. Here we plot only $n<1$. Error bars are estimated as the difference between densities with opposite $\alpha_{\mu}$ signs.
@@
To tie up all of these results, Figure 9 presents a heat map of $\partial c_{n}/\partial n$, marking the locations of the specific heat and density anomalies in the $n$ versus $U$ plane. Black diamonds indicate where the potential energy contribution of $c_{n}$, $c_{n,\text{pot}}$, is minimum, in connection with adiabatic cooling. As the repulsion increases, this minimum shifts towards densities closer to the half-filling. Black squares denote where $c_{n,\text{pot}}$ changes sign, pointing out the density values where double occupancy begins to be favored. As the interaction is increased, this curve crosses the $\alpha_{\mu}=0$ line at intermediate values of $U$, moving towards the half-filling. Filled black circles and crosses show, respectively, the loci of minimum values of $c_{n}$ and $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=\partial s/\partial n=0$. For $U\approx 6$ these crosses and circles are very close, as well as the change of sign in the potential energy contribution. Increasing the repulsion $U$ shifts the $c_{n}$ minimum to densities $n\approx 1$, while the zero crossing of $\alpha_{\mu}$ is shifted to densities away from half-filling. Furthermore, the region where $c_{n}$ displays local minima due to the strong correlations contains the curve $\alpha_{\mu}=0$. This can be understood as follows. The intensity of $\partial n_{\bm{k}}/\partial T$ for each $\bm{k}$ determines the kinetic contribution to $c_{n}$, while the relative values control the sign of $\alpha_{\mu}$. That is, a small kinetic energy contribution requires a small $\alpha_{\mu}$. This is why the $\alpha_{\mu}=0$ is and the region with $c_{n}$ minima in Figure 9 coincide.
@@
Figure 9: Heat map of $\left(\frac{\partial c_{n}}{\partial n}\right)_{T}$ for $T=1$ in the $n$ versus $U$ plane. Simulations were performed for $U\in\{2,4,6,7,8,9,10\}$ to produce an interpolation. Crosses mark states with $\alpha_{\mu}=0$. Filled circles denote where $c_{n}$ is minimum, filled squares represent $c_{n,\text{pot}}=0$ and black diamonds indicate the loci of minimum $c_{n,\text{pot}}$.
@@ IV Conclusions
Using Quantum Monte Carlo simulations, we studied the triple-peak structure of the specific heat as a function of the filling. This effect emerges in the low temperature regime $T\approx 1$ and for moderate to strong couplings $U\approx 6$, differently from the $c_{n}$ peak structure as a function of temperature, which exists even for weak interactions. As the filling of the lattice is increased below half-filling, a decreasing kinetic energy leads to a lower specific heat at constant density with increasing $U$, meaning it is easier to heat up the system for larger interaction strengths. This effect is explained by analyzing the momentum space distribution $n_{\bm{k}}$ of the interacting fermions, which is experimentally measurable in cold atom experiments. We detected an enhanced negative potential energy contribution at lower temperatures, which, through the double occupancy decrease with temperature, leads to adiabatic cooling at a wide range of densities. While the maxima threshold of $n=0.5$ in the specific heat does not seem to depend much on the interaction $U$ for the square lattice, it could be greatly affected by the lattice geometry. Future investigations could clarify the effects of frustration, disorder, or even long-range hopping on the described anomaly.
Connected to the specific heat through $\partial n_{\bm{k}}/\partial T$ we have shown the presence of a density anomaly, widely explored in the context of complex fluids and materials engineering. Its signature in the Fermi-Hubbard model is $\alpha_{\mu}<0$ below half-filling and $\alpha_{\mu}>0$ above it, which are found for strong correlations $U\gtrsim 6$. This density response also regulates thermoelectric properties: the sign of the Seebeck coefficient (in the long wavelength and static limits) is determined by the thermal expansion coefficient, as well as maxima in the entropy as a function of density. Therefore, Seebeck and density anomalies are tied together in this regime. Further research in this direction could analyze additional interaction scales, such as the extended Hubbard model, in analogy to classical systems with multiple potential wells or steps [55]. Understanding thermal anomalies is a crucial step to stretch the limits of cooling processes. For instance, the Mpemba effect, which is still the subject of debate in the case of water, is being studied with quantum simulators [56]: systems which are completely different are connected by how fast cooling or heating can happen.
Moreover, we note that a finite-temperature approximant for the Von Neumann entropy has been reported to show maxima as a function of density [57]. A possible avenue of future research is to understand how these maxima relate to the ones shown in our work.
@@ Acknowledgements.
We thank Natanael C. Costa for discussions about this work. We thank CENAPAD-SP (Centro Nacional de Processamento de Alto Desempenho em São Paulo), projeto UNICAMP/FINEP-MCTI for computational time for the DQMC calculations. M.A.H. thanks CAPES for funding and IIP-UFRN, where part of this work was completed (Grant No. 1699/24 IIF-FINEP). M.C.B. thanks CNPq for grants 403427/2021-5 and 311707/2021-1. Financial support from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, grant number E-26/204.308/2021 (W.C.F.S.), E-26/200.959/2022 (T.P.) and E-26/210.100/2023 (T.P.). T. P. also acknowledges support from CNPq grant numbers 308335/2019-8, 403130/2021-2, and 442072/2023-6.

### v2

@@ III Results
Because the denominator of Equation 9 is always positive due to thermodynamic stability in the single-band Hubbard model [57], a sign change of the Seebeck coefficient must be induced by a sign change of $\alpha_{\mu}$. Namely, the reported Seebeck anomaly [22] regarding its sign, and hence carrier charge, is controlled by the thermal expansion, with its magnitude modulated by the isothermal compressibility. This connection can be conceptually visualized as follows: a given temperature gradient $\delta T>0$ will induce a density change $\delta n>0$, as long as $\alpha_{\mu}$ is positive. This charge variation will induce an electric field in the same direction, and hence, we get a positive Seebeck coefficient. On the other hand, an opposite response of density $\delta n<0$, due to a negative $\alpha_{\mu}$, will establish an electric field in the opposite direction, yielding a negative Seebeck response. Near half-filling and for strong interactions, $n^{2}\kappa_{T}$ goes to zero [22], greatly enhancing the Seebeck coefficient. But the thermal expansion coefficient is zero at $n=1$ due to particle–hole symmetry, leading to a zero Seebeck coefficient. Microscopically, the change in carrier number above (below) half-filling is connected to the restructuring of the Fermi surface, which shifts from hole-like (electron-like) to electron-like (hole-like) above (below) half-filling [58].
Studies of classical lattice models connected a change of sign in $\alpha_{\mu}$ to residual entropies [16], as well as the results for bosons [14, 10]. Although the ground state of the Hubbard model has no residual entropy in the thermodynamic limit, at finite temperature, the entropy as a function of lattice filling $s(n)$ has maxima away from half-filling for strong correlations [22]. From Equation 9, we can see that $\alpha_{\mu}=0$ implies $(\partial s/\partial n)_{T}=0$. That is, the change of sign of $\alpha_{\mu}$ happens right at the entropy maxima. Conversely, the same is true for the Seebeck coefficient: its anomalous change of sign happens at the maxima in the entropy.
@@
To tie up all of these results, Figure 8 presents a heat map of $\partial c_{n}/\partial n$, marking the locations of the specific heat and density anomalies in the $n$ versus $U$ plane. Black diamonds indicate where the potential energy contribution of $c_{n}$, $c_{n,\text{pot}}$, is minimum, in connection with adiabatic cooling. As the repulsion increases, this minimum shifts towards densities closer to the half-filling. Black squares denote where $c_{n,\text{pot}}$ changes sign, pointing out the density values where double occupancy begins to be favored. As the interaction is increased, this curve crosses the $\alpha_{\mu}=0$ line at intermediate values of $U$, moving towards the half-filling. Filled black circles and crosses show, respectively, the loci of minimum values of $c_{n}$ and $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=\partial s/\partial n=0$. For $U\approx 6$ these crosses and circles are very close, as well as the change of sign in the potential energy contribution.
Increasing the repulsion $U$ shifts the $c_{n}$ minimum to densities $n\approx 1$, while the zero crossing of $\alpha_{\mu}$ is shifted to densities away from half-filling. Furthermore, the region where $c_{n}$ displays local minima due to the strong correlations contains the curve $\alpha_{\mu}=0$. This can be understood as follows. The intensity of $\partial n_{\bm{k}}/\partial T$ for each $\bm{k}$ determines the kinetic contribution to $c_{n}$, while the relative values control the sign of $\alpha_{\mu}$. That is, a small kinetic energy contribution requires a small $\alpha_{\mu}$. This is why the $\alpha_{\mu}=0$ is, and the region with $c_{n}$ minima in Figure 8 coincides. We show the entropy $s(n)$ and $\alpha_{\mu}=0$ data for other temperatures in Appendix C.
@@
Figure 8: Heat map of $\left(\frac{\partial c_{n}}{\partial n}\right)_{T}$ for $T=1$ in the $n$ versus $U$ plane. Simulations were performed for $U\in\{2,4,6,7,8,9,10\}$ to produce an interpolation. Crosses mark states with $\alpha_{\mu}=0$. Filled circles denote where $c_{n}$ is minimum, filled squares represent $c_{n,\text{pot}}=0$, and black diamonds indicate the loci of minimum $c_{n,\text{pot}}$. Results are for the $10\times 10$ lattice.
The kinetic part of the specific heat suggests carrier mobility influences the lateral maxima. Thus, we also analyzed the specific heat as a function of density when an additional hopping $t^{\prime}$ is added in the $\hat{x}+\hat{y}$ direction, breaking particle-hole symmetry. For $t^{\prime}=1$, the system is equivalent to a triangular lattice.
In Figure 9a), we show the cases $t^{\prime}=-0.2$, $t^{\prime}=0$ and also $t^{\prime}=1$. At half-filling, all of them show a peak, although the intensity is reduced at $t^{\prime}=1$ compared with the $t^{\prime}=0$ case. For larger doping, the loss of particle-hole symmetry is seen: the specific heat peak at $n\approx 1.5$ is larger for the $t^{\prime}=-0.2$ than the $t^{\prime}=0$ case, while the peak at $n\approx 0.5$ has the opposite behavior. For the $t^{\prime}=1$ case, the result is more intense, with the $n=1.5$ peak shifting to lower densities and greatly diminishing. Moreover, the peak at $n=0.5$ is significantly enhanced. For t=0.2t=-0.2, we observe the inverted behavior, but this specific value of hopping is too small for us to detect the lateral shift.
In panels b) and c) of Figure 9, we display the specific heat decomposition into kinetic and potential contributions with the new hopping term. Both the shift to lower densities and the intensity decrease in the case of $t^{\prime}=1$ are driven mostly by the kinetic contribution. In the potential term, we see a stronger $(\partial d/\partial T)_{n}<0$ below half-filling, but there is no adiabatic cooling above half-filling.
Figure 9: Specific heat decomposition for $U=10$, $T=1$ with an additional hopping $t^{\prime}$ in the $\hat{x}+\hat{y}$ direction in the $10\times 10$ lattice.
@@ IV Conclusions
Using Quantum Monte Carlo simulations, we studied the triple-peak structure of the specific heat as a function of the filling. This effect emerges in the temperature regime $J\lesssim T\ll U$, where $J=4t^{2}/U$. As the filling of the lattice is increased below half-filling, a decreasing kinetic energy leads to a lower specific heat at constant density with increasing $U$, meaning that it is easier to heat up the system for larger interaction strengths. The maximum threshold of $n=0.5$ in the specific heat does not depend much on the interaction $U$ for the square lattice. However, this was considerably altered in the triangular lattice, with differing heights and positions of maxima.
The reported effect complements the half-filling $c_{n}(T)$ known peak structure. We explained it by analyzing the momentum space distribution $n_{\bm{k}}$ of the interacting fermions, showing the low–$k$ electrons become less sensitive to temperature at the interactions and lattice fillings where $c_{n}(n,U)$ behaves differently from the non-interacting case. The distribution $n_{\bm{k}}$ is experimentally measurable in cold atom experiments. Moreover, we detected an enhanced negative potential energy contribution at lower temperatures. Due to double occupancy decreasing with temperature, this leads to adiabatic cooling at a wide range of densities.
Connected to the specific heat through $\partial n_{\bm{k}}/\partial T$ we have shown the presence of a density anomaly, widely explored in the context of complex fluids and materials engineering. Its signature in the Hubbard model is $\alpha_{\mu}<0$ below half-filling and $\alpha_{\mu}>0$ above it, which are found for strong correlations $U\gtrsim 6$. This density response also regulates thermoelectric properties: the sign of the Seebeck coefficient (in the long wavelength and static limits) is determined by the thermal expansion coefficient, as well as maxima in the entropy as a function of density. Therefore, Seebeck and density anomalies are tied together in this regime. Further research in this direction could analyze additional interaction scales, such as the extended Hubbard model, in analogy to classical systems with multiple potential wells or steps [59].
Understanding thermal anomalies is a crucial step to stretching the limits of cooling processes. For instance, the Mpemba effect, which is still the subject of debate in the case of water, is being studied with quantum simulators [60]: systems which are completely different are connected by how fast cooling or heating can happen. Furthermore, we note that a finite-temperature approximant for the Von Neumann entropy has been reported to show maxima as a function of density [61]. A possible avenue of future research is to understand how these maxima relate to the ones shown in our work.
@@ Acknowledgements.
We thank Natanael C. Costa and Rodrigo G. Pereira for discussions about this work. We thank CENAPAD-SP (Centro Nacional de Processamento de Alto Desempenho em São Paulo), projeto UNICAMP/FINEP-MCTI for computational time for the DQMC calculations. M.A.H. thanks CAPES for funding and IIP-UFRN, where part of this work was completed (Grant No. 1699/24 IIF-FINEP). M.C.B. thanks CNPq for grants 403427/2021-5 and 311707/2021-1. Financial support from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, grant number E-26/204.308/2021 (W.C.F.S.), E-26/200.959/2022 (T.P.), E-26/210.100/2023 (T.P.), and E-26/210.781/2025 (T.P.), T. P. also acknowledges support from CNPq grant numbers 308335/2019-8, 403130/2021-2, and 442072/2023-6.
Appendix A $c_{n}$ derivation
The specific heat at constant density is defined as
$c_{n}=\left(\frac{\partial E}{\partial T}\right)_{n}\;.$ (10)
Since we are working in the Grand Canonical Ensemble, $E=E(T,n)=E(T,n(T,\mu))$, hence
$\left(\frac{\partial E}{\partial T}\right)_{n}=\left(\frac{\partial E}{\partial T}\right)_{\mu}+\left(\frac{\partial E}{\partial\mu}\right)_{T}\left(\frac{\partial\mu}{\partial T}\right)_{n}\;.$ (11)
We use that
$\left(\frac{\partial\mu}{\partial T}\right)_{n}=-\frac{\left(\frac{\partial n}{\partial T}\right)_{\mu}}{\left(\frac{\partial n}{\partial\mu}\right)_{T}}\;.$ (12)
Hence,
$\left(\frac{\partial E}{\partial T}\right)_{n}=\left(\frac{\partial E}{\partial T}\right)_{\mu}-\frac{\left(\frac{\partial E}{\partial\mu}\right)_{T}\left(\frac{\partial n}{\partial T}\right)_{\mu}}{\left(\frac{\partial n}{\partial\mu}\right)_{T}}\;.$ (13)
Identifying
$c_{\mu}$ $=\left(\frac{\partial E}{\partial T}\right)_{\mu}\;,$ (14)
$\alpha_{\mu}$ $=\left(\frac{\partial n}{\partial T}\right)_{\mu}\;,$ (15)
$\kappa_{T}$ $=\frac{1}{n^{2}}\left(\frac{\partial n}{\partial\mu}\right)_{T}\;,$ (16)
we arrive at Equation 2.
Appendix B Sign problem & error propagation
The DQMC technique is unbiased but suffers from the well-known sign problem [33, 32]. This requires a reweighting of the observables:
$\braket{A}=\frac{\braket{\text{sgn}\times A}}{\braket{\text{sgn}}}.$ (17)
That is, instead of measuring the observable $A$, we measure $\text{sgn}\times A$ and normalize by the average sign. If the average sign $\braket{\text{sgn}}$ is close to zero, good statistical averages for the observable require exponentially longer simulations. The repulsive Hubbard model is sign-problem-free at half-filling ($\mu=0$), but doping changes this scenario. In general, the sign problem becomes worse at lower temperatures, in larger systems, and with stronger interactions. In our work, we restricted ourselves to temperatures at which the sign problem was not severe. We show out lowest temperature $T=0.5$ in Figure B1 for the $10\times 10$ lattice.
Figure B1: Average sign at $T=0.5$ for various interactions $U$. Points are for the $N=10\times 10$ lattice unless the label specifies otherwise. The black points correspond to the calculated fillings for the largest lattice shown in Figure 3. Continuous lines are a guide to the eyes.
The DQMC simulations measures $\braket{A}$ multiple times, leading to an average $\overline{A}$ and standard deviation $\delta A$. We can then calculate numerical derivatives through finite differences, taking into account the error bars. For instance, the temperature derivative of the internal energy and its error are calculated with
$\left(\frac{\partial E}{\partial T}\right)_{\mu}$ $=-\beta^{2}\frac{E(\beta+2\Delta\tau,\mu)-E(n-2\Delta\tau,\mu)}{4\Delta\tau}$
$\delta\left(\frac{\partial E}{\partial T}\right)_{\mu}$ $=\beta^{2}\frac{\sqrt{\delta E(\beta+2\Delta\tau,\mu)^{2}+\delta E(n-2\Delta\tau,\mu)^{2}}}{4\Delta\tau}$
Note that in DQMC there is an imaginary time dimension which is discretized in $\Delta\tau$ steps. Hence the finite difference must be applied in temperatures which are an integer multiple of $\Delta\tau$. In particular, for $T=0.5$ we used $\Delta\tau=0.05$ and $L=36,40,44$.
Appendix C Entropy for various temperatures
Figure C2 shows that we detect the known maxima away from half-filling for all simulated temperatures. While for $T=2$ only $U=8$ and $U=10$ show maxima away from half-filling, at $T=0.5$ this effect begins already at $U=4$. At the extreme values of $n=0$ and $n=2$ the system is a band insulator, leading to zero entropy. Note that at the low temperature the $U=10$ case has the highest entropy and $U=0$ the lowest, while the opposite happens at $T=2$, for all fillings. The $T=0.5$ results were shown previously in [22].
Figure C2: Entropy $s$ as a function of density $n$ for a) $T=2$, b) $T=1$ and c) $T=0.5$ considering interactions ranging from the weak to the strong coupling limit. The $U=0$ line is hidden behind the $U=2$ case.
To map the location of the peaks as temperature changes, we mark on Figure C3 the points where $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=\partial s/\partial n=0$, for all simulated temperatures and interactions. For the higher temperature, strong interactions ($U>8$) are necessary for the change of sign to occur. As the temperature is lowered, the effect starts to occur at a lower value of $U$. Our results are in quantitative agreement with behavior of the peaks available in the literature, even thought the previous calculation was for a smaller $4\times 4$ system [28] which resulted in additional entropy peaks at other fillings.
Figure C3: Location of $\alpha_{\mu}=\mathcal{S}_{\text{Kelvin}}=(\partial s/\partial n)_{T}=0$ for all the simulated temperatures and interactions, estimated by the average between densities with opposite $\alpha_{\mu}$ signs. Here we plot only $n<1$. Error bars are estimated as the difference between densities with opposite $\alpha_{\mu}$ signs.
Appendix D Comparison of specific heat derivative at constant density and chemical potential
We show in Figure D4 the momentum-space temperature derivative of the occupation $n_{\bm{k}}$ at constant density $n$. In a similar manner as shown for the constant-$\mu$ derivative, the strong interaction at $n\approx 0.83$ leads to a low temperature sensitivity in comparison with the other panels.
Figure D4: The momentum distribution of $\left(\frac{\partial n_{\bm{k}}}{\partial T}\right)_{n}$ in the first Brillouin zone for interactions $U=2$ (upper panels) and $U=10$ (lower panels), and fillings $n\approx 0.3$ (left panels), $n\approx 0.51$ (central panels) and $n\approx 0.83$ (right panels). The temperature is $T=1$ and the lattice size is $24\times 24$, with a cubic spline interpolation between points.